the first side of a tringle is 8cm shorter than the second side,the third side is 4 times as long as the first side. the perimiter is 26cm. find the lenght of the sides.
Answer by Paul(988)

Second side = x
First side = x-8
THird side = 4(x-8)
P=26
Together:

x+(x-8)+4(x-8)=26
2x-8+4x-32=26
6x-40=26
6x=66
x=11 (second sidde)

11-8= 3(first side)

4(11-3)=12 (third side)

Hence, the three sides are 3, 11, 12.
Paul.